OpenStudy (anonymous):

HELP! Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (4, -5).

OpenStudy (anonymous):

slope-intercept form is y=mx+c m=gradient or also known as the slope c=y-intercept so we use m=-3 and the point (4,-5) and substitute it to find the value of c. then just re-write the whole equation with the gradient and c.

OpenStudy (anonymous):

or another method is using the find an unknown equation formula. y-y1 = m(x-x1) substitute the point (4,-5) as (x1,y1) and also m as the gradient then rearrange into the form y = mx + c